# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v3

@@ 5 Experiments
We now present empirical results to answer the questions:
  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR?
  • What about the generalization ability of TAD in both image-based and state-based environments?
  • Can TAD be extended to more general settings like dynamic generalization?
@@ 5.1 Experimental Setup
To answer the first question and verify analyses about TDR, we first consider lots of different task combinations in DeepMind control suite (DMC) Tassa et al. (2018). These task combinations include (1) two task combinations of Walker-stand&\&&walk Mu et al. (2022) and Acrobot-swingup&\&&sparse, which share the same optimal actions with TDR being 0; (2) two task combinations with non-zero TDR of Reacher-easy&\&&hard Mu et al. (2022) and Walker-walk&\&&prostrate. For example, Walker-walk and Walker-prostrate hope a two-leg robot to maintain its height above a certain threshold for walking and below another certain threshold respectively, yielding almost opposite optimal Q functions with huge TDR. In Appendix C.1, we introduce more details about them and four other task combinations.
To evaluate generalization, we extend tasks in DMC and design three task distributions: (1) Cheetah_normal-_\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum_normal-_\__angle($\alpha,\ \beta$), which extends Pendulum-swingup and hopes an inverted pendulum to stay its pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker_normal-_\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$. For each task distribution, we sample $4$ tasks for training and $2$ additional tasks for testing. More details are in Appendix C.2.
@@
To demonstrate the scalability of TAD, we also consider some state-based continuous control task distributions about robotic locomotion, simulated via the MuJoCo simulator Todorov et al. (2012). Following previous work Finn et al. (2017), our chosen tasks distributions include (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 testing tasks. More details of these task distributions are in Appendix D.
@@
In DMC, we choose two model-free methods: CURL Laskin et al. (2020) and SAC+AE Yarats et al. (2021), which employ Markovian policies and belong to $\Pi_{1}$. In addition, we choose two classic world models, PlaNet Hafner et al. (2019b) and Dreamer Hafner et al. (2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$. In state-based control, besides PlaNet and Dreamer, we also take some meta RL methods like MAML Finn et al. (2017), RL2 Duan et al. (2016), and VariBAD Zintgraf et al. (2019), including zero-shot and few-shot evaluation, as reference.
@@
For the task combinations, we evaluate the average return of all tasks to verify TDR. For the task generalization setting, we train agents in training tasks and evaluate their generalization abilities in testing tasks. For all experiments, we train 3 policies with different seeds and calculate the mean and std to mitigate the effects of randomness.
Train&\&&Test Train&\&&Test Train Test Train Test
@@
Figure 2: Visualization of Dreamer (upper) and TAD (below) for handling Walker-walk&\&&Prostrate. For each algorithm in each task, we show 4 frame observations of a trajectory in the right part. In the left part, we draw t-SNE results of all states in trajectories of Walker-walk (blue points) and Walker-prostrate (red points).
5.2 Experimental Results for TDR
To answer the first question and validate our analyses of TDR, we report the optimal mean $\pm$ std of the average cumulative return of different task combinations in Table 5. Comprehensive results of all 8 task combinations and training curves are in Appendix C.3. The results show that TAD outperforms baselines in all environments, especially in those with high TDR. In task combinations where TDR is 0, like Walker-stand&\&&walk or Acrobot-swingup&\&&sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, as derived in Theorem 3. In these cases, $\Pi_{1},\Pi_{2}$ performs well and TAD can surpass baselines in Acrobot-swingup&\&&sparse since it can better distinguish different tasks. In the latter two environments with non-zero TDR, TAD achieves a notable improvement compared with baselines. Especially, in Walker-walk&\&&prostrate, where two tasks own almost opposite optimal Q functions and TDR is significantly huge, policies belong to $\Pi_{1},\Pi_{2}$ and can not differentiate different tasks, aligning with our Theorem 3. In contrast, TAD can distinguish between tasks by utilizing $\Pi_{3}$ and improves performance conspicuously.
Moreover, we present some visualization to better understand how TAD works in Fig. 2. We sample trajectories from the trained agent via Dreamer as well as TAD in Walker-walk&\&&prostrate and visualize some observations (right part of Fig. 2) as well as all states (left part of Fig. 2), of which the dimensions are reduced for visualization by t-SNE Van der Maaten and Hinton (2008). As shown here, states in TAD of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks. Consequently, TAD can successfully distinguish different tasks and handle them at the same time effectively. Videos of these trajectories are in supplementary materials.
5.3 Experimental Results for Task Generalization
To answer the second question, we report the results under the task generalization setting of image-based and state-based environments in Table 2 and Table 3, respectively. In Table 2, TAD shows powerful generalization ability in image-based environments as it can significantly enhance the average cumulative return in both training tasks and testing tasks compared with baselines. This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen testing tasks effectively. We also provide some visualization results in Appendix C.4 with corresponding videos in supplementary materials.
Additionally, in Table 3, we compare TAD with some meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively. Many context-based meta-RL methods, including RL2, PEARL and VariBAD, utilize reward into policies and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses. As shown in Table 3, TAD can achieve a significant improvement in the average cumulative return for both training tasks and testing tasks, since TAD is aware of task information and can generalize to unseen tasks.
5.4 Ablation Study
Reward Signal and Task ID.
In Fig. 3, we do ablation studies with two baselines: Dreamer(w/ r), which only integrates $\Pi_{3}$ into Dreamer, i.e., including reward signals into policies, and Dreamer(w/ task ID), which includes one-hot task ID into policies in both training and testing. Results show that just providing reward signals can help Dreamer distinguish different tasks, which verifies our analyses in TDR. Also, TAD shows superior performance even compared with Dreamer(w/ task ID), which demonstrates that our task model and corresponding ELBO are effective for task generalization.
Figure 3: Ablation study on Reward Signal and Task ID.
Extension to Dynamic Generalization.
To answer the third question, we further evaluate TAD in more general settings with different observations, dynamics, and/or actions. As TAD utilizes all historical information to infer the environment, it can be directly applied to these settings. We design task distribution with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run__\__mass, Walker-walk__\__mass). Table 4 shows that TAD can achieve much greater performance and better convergence compared to Dreamer, which indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks. More details about the environments and results are in Appendix E.1.
5.5 Limitations and Ethical Statement
In terms of limitations, TAD assumes that the task context is continuously related to historical information for generalizing to unseen tasks. Thus TAD might be difficult to generalize in sparse-reward settings. We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix B.7). Fortunately, in relatively sparse reward settings, we conduct studies in Appendix E.2 to show that TAD can infer the current task and generalize to unseen tasks well. There are no serious ethical issues as this is basic research.
6 Conclusion
In this work, we propose a novel framework of TAD that distinguishes different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features. In TAD, we calculate the corresponding variational lower bound of the data log-likelihood, which includes a novel loss term to distinguish different tasks via states. To explain components in TAD, we introduce a novel metric TDR to capture the relevance of the task distribution and show that Markovian policies perform poorly in tasks with high TDR. Experiments in image-based and state-based settings demonstrate that TAD can remarkably improve the performance of handling different tasks meanwhile, especially for high TDR ones, and successfully generalize to unseen tasks.
Appendix A Pseudo Code of TAD
We provide the pseudo code of TAD in Algorithm 2.
@@ Appendix A Pseudo Code of TAD
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
@@
Appendix B Proof of Theorems
In this section, we will provide detailed proofs of theorems.
B.1 The Proof of Theorem 1
In this part, we give an informed version of Theorem 1 with detailed proof as
Theorem 5.
Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, we consider the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M state-Q function spaces, $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is some Q function. Given $M$ corresponding dataset $D^{m}=\{(s_{t}^{m},a_{t}^{m},r_{t}^{m},s_{t+1}^{m})\}$, we can consider three hypothesis classes of optimal Q function $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$, which are subsets of $\mathcal{H}$:
$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(s_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(s_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall s,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p(s,a),a^{\prime}),\forall s,a\}.\end{split}$ (8)
Proof.
First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in Young et al. (2022) that considers the setting that the rewards of all environments are the same.
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists\{p_{m}\}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall m,s,a\\ \Rightarrow&\exists\{p_{m}\}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(s_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (9)
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p(s,a),a^{\prime}),\forall m,s,a\\ \Rightarrow&\exists p_{m}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall m,s,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (10)
thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
B.2 The Derivation of the ELBO
Thus we have proven the ELBO.
B.3 The Proof of Theorem 2
Proof.
Thus we have proven this result. ∎
B.4 The Proof of Theorem 3
Proof.
Our proof follows some previous work Kakade and Langford (2002); Ying et al. (2022). First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
here $\Delta\pi(a|s)=\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)$. Since Eq. (14) holds for any $s$, thus we calculate its expectation for $s\sim d_{\mathcal{M}}^{\pi_{\mathcal{M}}^{*}}$:
thus we can calculate that
[…]
Consequently, we have
Thus we have proven this result. ∎
B.5 Expressive Ability of $\Pi_{3}$
Proposition 1.
Proof.
here $i=\arg\max\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{n-1}\}+1$, thus we have
Thus we have proven this result. ∎
B.6 Proof and Discussion of Theorem 4
Theorem 6.
For any policy $\pi\in\Pi_{3}$, we have
Proof.
$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq. (18), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same. Thus we need to change the Eq. (19) as
[…]
Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization. In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution. In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by E.q. (27) will be controlled.
B.7 Sparse Reward Task
$\mathcal{R}_{\mathcal{M}_{i}}(s_{T-1},a_{i})=1,i=1,...,n.$ (29)

### v4

@@ 5 Experiments
We now present empirical results to answer the following questions:
  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR? (Sec. 5.2)
  • How about TAD’s generalization ability when handling image-based and state-based observations? (Sec. 5.3)
  • Can TAD be extended to more general settings like dynamic generalization? (Sec. 5.4)
@@ 5.1 Experimental Setup
To verify our analyses of TDR in Sec. 4.3, we consider several task combinations in DeepMind Control suite (DMC) (Tassa et al. 2018): (1) Cartpole-balance&balance__\__sparse, which shares the same optimal actions with TDR 0; (2) two task combinations with non-zero TDR of Walker-stand&walk&prostrat&flip, and Cheetah-run&run__\__back&flip&flip__\__flip__\__back, which are widely used in multi-task unsupervised RL (Sekar et al. 2020; Laskin et al. 2021). For example, Cheetah-run and Cheetah-flip hope a two-leg robot to move forward by running and flip around the torso, respectively (see Fig. 3), yielding almost opposite optimal Q functions with huge TDR. In Appendix C.1, we introduce more details about these task combinations.
To evaluate generalization with image-based observations, we extend tasks in DMC and design three task distributions: (1) Cheetah__\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum__\__angle($\alpha,\ \beta$), extending Pendulum-swingup to keep pendulum’s pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker__\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$. For each task distribution, we sample $4$ training tasks and $2$ additional test tasks. More details are in Appendix C.2.
@@
To demonstrate the scalability of TAD, we also consider some state-based continuous robotic control task distributions simulated via MuJoCo (Todorov, Erez, and Tassa 2012). Following previous work (Finn, Abbeel, and Levine 2017), we choose tasks distributions: (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 test tasks. More details of these task distributions are in Appendix D.
@@
In DMC, we choose two model-free methods employing Markovian policies $\Pi_{1}$: CURL (Laskin, Srinivas, and Abbeel 2020) and SAC+AE (Yarats et al. 2021). In addition, we choose two classic world models, PlaNet (Hafner et al. 2019b) and Dreamer (Hafner et al. 2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$. Moreover, we take a SOTA model-based meta RL method belonging to $\Pi_{3}$: MAMBA (Rimon et al. 2024). In state-based control, besides PlaNet and Dreamer, we take some meta RL methods like MAML (Finn, Abbeel, and Levine 2017), RL2 (Duan et al. 2016), and VariBAD (Zintgraf et al. 2019), including zero-shot and few-shot evaluation, as reference.
@@
For the task combinations, we evaluate the average return of all tasks to verify TDR. For the task generalization settings, we train agents in training tasks and evaluate their generalization abilities in test tasks. For all experiments, we repeat 5 different random seeds and report the mean $\pm$ std to mitigate the effects of randomness following previous works (Hafner et al. 2019a; Rimon et al. 2024).
Figure 3: Sampled trajectories and imaginary trajectories of TAD for different tasks (Cheetah-run and Cheetah-flip).
Figure 4: The t-SNE clustering of state embeddings for different tasks sampled via Dreamer, TAD-CE, and TAD-SC.
5.2 Experimental Results for TDR
To validate our analyses of TDR, we report different algorithms’ performance of different task combinations in Table 1. Results show that TAD-CE and TAD-SC outperform all baselines, especially in those environments with high TDR. As derived in Theorem 3, in task combinations where TDR is 0 like Cartpole-balance&balance__\__sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, thus baselines like CURL and Dreamer perform well. Conversely, in other task combinations, different tasks’ optimal Q functions may differ a lot (like Chetah-run and Cheetah-run__\__back) and TDR is significantly huge. Thus methods with policies in $\Pi_{1},\Pi_{2}$ like CURL and Dreamer can not differentiate different tasks and perform poorly. Aligning with Theorem 3, baselines utilizing $\Pi_{3}$ like MAMBA outperform other baselines and TAD improves performance conspicuously.
Moreover, we present some visualization to better understand how TAD works. Fig, 3 shows video predictions of TAD for different tasks (Cheetah-run and Cheetah-flip). We use the same agent trained by TAD to sample trajectories for these two tasks and show their trajectories in lines 1 and 3 respectively. Given the first 7 steps as the context (their observations are similar but the rewards are different), we directly imagine the future 55 steps by our trained reward-informed world model, of which the results are in lines 2 and 4 of Fig. 3 respectively. As shown here, TAD performs differently to handle different tasks only by receiving different rewards. Moreover, our reward-informed world models demonstrate long-term prediction and high-quality reconstruction capabilities, which reveals the potential for training future large-scale world models. In Fig. 4, we sample trajectories from agents trained by Dreamer, TAD-CE, as well as TAD-SC, and visualize the states of different tasks, of which the dimensions are reduced for visualization by t-SNE (Van der Maaten and Hinton 2008). As shown here, states in TAD-CE and TAD-SC of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks. This result demonstrates TAD can effectively learn task-aware information and distinguish different tasks. Videos of different tasks are provided in supplementary materials.
5.3 Experimental Results for Task Generalization
To answer the second question, we report the task generalization results of image-based and state-based environments in Table 2-3. In Table 2, TAD shows more powerful generalization abilities than all baselines. This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen test tasks effectively. We also provide some visualization results in Appendix C.3 with videos in supplementary materials.
Additionally, in Table 3, we compare TAD with existing meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively. Many context-based meta RL methods, including RL2, PEARL, and VariBAD, utilize historical rewards and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses. As shown in Table 3, TAD achieves a significant improvement for both training tasks and test tasks, since TAD is aware of task information for generalizing to unseen tasks.
5.4 Ablation Study
Reward Signals.
In Fig. 5, we do ablation studies about the task term in TAD, i.e., we consider Dreamer(w/ r) that only integrates $\Pi_{3}$ into Dreamer. Results show that just providing rewards can help Dreamer distinguish different tasks, which verifies our analyses in TDR. Also, TAD-CE and TAD-SC show superior performance, demonstrating that our task model and corresponding ELBO are effective for task generalization. Also, we do ablation studies on reward signals for model-free methods like CURL in Appendix E.1.
Figure 5: Ablation study on Reward Signal.
Extension to Dynamic Generalization.
To answer the third question, we evaluate TAD in more general settings with different observations, dynamics, and/or actions. As TAD utilizes all historical information to infer the environment, it can be directly applied to these settings. We design task distribution with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run__\__mass, Walker-walk__\__mass). More details about environments and results are in Appendix E.2, where TAD achieves much greater performance compared to baselines and show potential in further handling dynamic generalization and even cross-embodiment tasks.
5.5 Limitations and Discussion
In terms of limitations, TAD assumes that the task context is continuously related to historical information for generalizing to unseen tasks. Thus TAD might be difficult to generalize in sparse-reward settings. We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix A.7). Fortunately, in relatively sparse reward settings, we conduct experiments in Appendix E.3 to show that TAD can infer the current task and generalize to unseen tasks well.
6 Conclusion
In this work, we propose a novel framework of TAD that handles different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features. In TAD, we calculate the corresponding variational lower bound of the data log-likelihood, which includes a novel loss term to distinguish different tasks via states. To explain components in TAD, we introduce a novel metric TDR to capture the relevance of the task distribution and show that Markovian policies perform poorly in tasks with high TDR. Experiments in image-based and state-based settings demonstrate that TAD can remarkably improve the performance of handling different tasks meanwhile, especially for high TDR ones, and successfully generalize to unseen tasks.
Appendix A Proof of Theorems
In this section, we will provide detailed proofs of theorems.
A.1 The Proof of Theorem 1
Proof.
First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in (Young et al. 2023) that considers the setting that the rewards of all environments are the same.
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\\ \Rightarrow&\exists(r_{m})_{m=1}^{M}::p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (10)
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall m,o,a\\ \Rightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall m,o,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (11)
thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
A.2 The Derivation of the ELBO
Thus we have proven the ELBO.
A.3 The Proof of Theorem 2
Proof.
Thus we have proven this result. ∎
A.4 The Proof of Theorem 3
Proof.
Our proof follows some previous work (Kakade and Langford 2002; Ying et al. 2022). First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
here $\Delta\pi(a|s)=\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)$. Since Eq. (LABEL:15) holds for any $s$, thus we calculate its expectation for $s\sim d_{\mathcal{M}}^{\pi_{\mathcal{M}}^{*}}$:
thus we can calculate that
@@ A.4 The Proof of Theorem 3
Consequently, we have
Thus we have proven this result. ∎
A.5 Expressive Ability of $\Pi_{3}$
Proposition 1.
Proof.
here $i=\arg\max\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{n-1}\}+1$, thus we have
Thus we have proven this result. ∎
A.6 Proof and Discussion of Theorem 4
Theorem 5.
For any policy $\pi\in\Pi_{3}$, we have
Proof.
$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq. (LABEL:app_eq_11), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same. Thus we need to change the Eq. (LABEL:app_eq_12) as
[…]
Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization. In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution. In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by Eq. (LABEL:app_ep_20) will be controlled.
A.7 Sparse Reward Task
$\mathcal{R}_{\mathcal{M}_{i}}(s_{T-1},a_{i})=1,i=1,...,n.$ (30)
Appendix B Pseudo Code of TAD
The detailed pseudo code of TAD is provided in Algorithm 2.
@@ Appendix B Pseudo Code of TAD
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic